OpenStudy (anonymous):

4(square root of seven)+3(square root of 7)

OpenStudy (anonymous):

\[4\sqrt{7}+3\sqrt{7}\]Pretend that the sqrt(7)s are variables.

OpenStudy (anonymous):

\[7\sqrt{7}?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yay! Thank you

OpenStudy (anonymous):

No problem